# Mixed Membership Estimation for Weighted Networks

### v2

@@ 5 Estimation of the Number of Communities
In this section, we aim at answering Q (3). Algorithm 1 requires a known number of communities $K$. However, for real-world networks, $K$ is usually unknown, which diminishes the usefulness of a community detection approach. Here, to estimate $K$, we introduce overlapping weighted modularity.
@@
$Q^{+}=\frac{1_{m^{+}>0}}{m^{+}}\sum_{k=1}^{K}\sum_{i\in\hat{\mathcal{C}}_{k},j\in\hat{\mathcal{C}}_{k}}\frac{1}{O(i)O(j)}(A^{+}(i,j)-\frac{d^{+}(i)d^{+}(j)}{m^{+}}),Q^{-}=\frac{1_{m^{-}>0}}{m^{-}}\sum_{k=1}^{K}\sum_{i\in\hat{\mathcal{C}}_{k},j\in\hat{\mathcal{C}}_{k}}\frac{1}{O(i)O(j)}(A^{-}(i,j)-\frac{d^{-}(i)d^{-}(j)}{m^{-}}),$
where $1_{m^{+}>0}$ is indicator function such that $11_{m^{+}>0}=$ 1 if $m^{+}>0$ and 0 otherwise. Similar for $1_{m^{-}>0}$. Following the idea of defining weighted modularity for signed networks gomez2009analysis , our overlapping weighted modularity is defined as
$Q=\frac{m^{+}}{m^{+}+m^{-}}Q^{+}-\frac{m^{-}}{m^{+}+m^{-}}Q^{-}.$ (7)
@@
Our overlapping weighted modularity evaluates the quality of overlapping community partition. Similar to the Newman-Girvan modularity newman2004finding ; newman2006modularity and weighted modularity gomez2009analysis , a larger overlapping weighted modularity indicates a better community partition. Our overlapping weighted modularity includes several popular modularity as special cases.
  • 1.
@@
    When $A$ has both positive and negative entries and $\hat{\Pi}^{0}$ degenerates such that each node only belongs to one community, our overlapping weighted modularity reduces to the modularity for signed networks provided in Equation (17) of gomez2009analysis .
  • 2.
@@
    When all entries of $A$ are nonnegative (i.e., $m^{-}=0$), our overlapping weighted modularity degenerates to the overlapping modularity provided in Equation (2) of shen2009detect .
  • 3.
@@
    When $\hat{\Pi}^{0}$ degenerates such that each node only belongs to one community and all entries of $A$ are nonnegative, our overlapping weighted modularity degenerates to the famous Newman-Girvan modularity newman2004finding ; newman2006modularity .
To infer the number of communities, we adopt the suggestion proposed in nepusz2008fuzzy , that is, we iteratively increase $k$ and choose the one maximizing our overlapping weighted modularity in Equation (7). We let $\hat{\Pi}^{0}$ be obtained from our ScD algorithm and call our method for inferring $K$ as KScD.
@@ 6 Experiments
We conduct extensive experiments to validate our theoretical results, demonstrate the effectiveness of ScD in estimating membership matrix $\Pi$, and demonstrate the effectiveness of our overlapping weighted modularity by showing that KScD can infer $K$ for networks generated from DCMMDF.
@@ 6.1 Baseline Methods
For the task of estimating memberships, we select three mixed membership community detection approaches as baseline methods of our DCMMDF model.
  • 1.
    GeoNMF GeoNMF adapts nonnegative matrix factorization to estimate mixed memberships for networks generated from MMSB.
  • 2.
    SPACL mao2020estimating is a spectral algorithm for estimating mixed memberships for networks generated from MMSB.
  • 3.
    Mixed-SCORE MixedSCORE is a spectral algorithm for estimating mixed memberships for networks generated from DCMM.
@@
GeoNMF and Mixed-SCORE may fail to output for the case that some elements in the adjacency matrix $A$ are negative. To make GeoNMF and Mixed-SCORE work when $A$ has negative elements, we make all entries of $A$ nonnegative by adding a sufficiently large constant for these two methods.
@@
  • 3.
    NB and BHac le2022estimating are two methods for inferring the number of communities and they determine $K$ via non-backtracking matrix and Bethe Hessian matrix, respectively.
Remark 5.
For determining $K$, we do not consider KGeoNMF which is defined similarly to KScD, KSPACL, and KMixed-SCORE because KGeoNMF fails to output for large $k$.
@@ 6.2 Evaluation metrics
Metrics with ground truth. For this case, we introduce three widely used evaluation metrics: Hamming error MixedSCORE and Relative error mao2020estimating .
  • 1.
@@
Metrics without ground truth. For this case, our overlapping weighted modularity in Equation (7) is capable to evaluate the quality of detected communities, and a higher value of overlapping weighted modularity indicates more accurately detected communities.
Accuracy rate. For the task of estimating $K$, we use the Accuracy rate to evaluate the precision of KScD and its competitors, where Accuracy rate is defined as the proportion of times an approach exactly determines $K$ le2022estimating .
@@ 6.3 Simulations
For computer-generated networks, we set $n=400$ and $K=3$. Let each community has $n_{0}$ number of pure nodes. For pure node $i$, let $\Pi(i,k)=1$ if $1+(k-1)n_{0}\leq i\leq kn_{0}$ for $i\in[Kn_{0}],k\in[K]$. Set $n_{0}=40$. Let all mixed nodes have four different memberships $(0.1,0.1,0.8),(0.1,0.8,0.1),(0.8,0.1,0.1)$, and $(1/3,1/3,1/3)$, each with $\frac{n-3n_{0}}{4}$ number of nodes. Let $P$’s diagonal elements be 1 and non-diagonal elements be $p$. For Normal distribution and signed networks, $p$ is set to be negative. For other distributions, $p$ is set to be positive. Let $\theta(i)=\rho\times(\mathrm{rand}(1)/2+0.5)$, where $\mathrm{rand}(1)$ is a random value ranged in $(0,1)$. In this way, $n,K,P,\Pi$, and $\Theta$ have been set. Then following Steps (a)-(c), we can generate an adjacency matrix $A$ with true membership matrix $\Pi$ for any distribution $\mathcal{F}$ satisfying Equation (6) under our DCMMDF model. For the task of community detection, applying ScD (and its competitors) to $A$ with $K$ communities obtains the estimated membership matrix $\hat{\Pi}$, and then we can obtain Hamming error and Relative error. For each parameter setting considered in this section, we report the averaged Hamming error (and Relative error) over 50 repetitions for each method. For the task of determining the number of communities, applying KScD (and its competitors) to $A$ obtains the estimated number of communities, and then we can obtain the Accuracy rate over the 50 repetitions.
@@
For our synthetic networks, $P,\Pi$, and $\Theta$ can be set arbitrarily as long as Equations (2)-(5) hold and $P$’s non-diagonal elements should be positive or nonnegative or can be negative depending on distribution $\mathcal{F}$ just as analyzed in Examples 1-5. Meanwhile, for simulations, $\mathcal{F}$ can be any distribution since our DCMMDF is a distribution-free model and it has no constraint on a specific distribution $\mathcal{F}$ as long as Equation (6) holds. Furthermore, to save space, we do not consider missing edges in the simulations.
Experiment 1: Normal distribution. Set $\mathcal{F}$ as Normal distribution such that $A(i,j)\sim\mathrm{Normal}(\Omega(i,j),\sigma^{2}_{A})$. Let $p=-0.2,\sigma^{2}_{A}=2$, and $\rho$ range in $\{0.1,0.2,\ldots,2\}$. The results are in panels (a)-(c) of Figure 3. We observe the following: (i) ScD performs better as $\rho$ increases, and this is consistent with our theoretical analysis in Example 1. (ii) ScD, GeoNMF, SPACL, and Mixed-SCORE perform similarly and they all successfully estimate memberships for large $\rho$. (iii) For the task of inferring the number of communities, our methods KScD, KSPACL, and KMixed-SCORE designed based on our overlapping weighted modularity enjoy similar and satisfactory performances because their accuracy rates increase as $\rho$ increases, which indicates the effectiveness of our overlapping weighted modularity. For comparison, NB and BHac fail to estimate $K$ for Normal distribution.
@@
For the task of estimating memberships, in our experimental studies, we aim at using simulations to verify our theoretical analysis that ScD has different behaviors for different distributions when $\rho$ is changed as analyzed in Examples 1-5. We also want to investigate the performances of GeoNMF, SPACL, and Mixed-SCORE when they are applied to estimate memberships for weighted networks generated from our model DCMMDF instead of showing that ScD outperforms its competitors.
@@
Experiment 2: Bernoulli distribution. Set $\mathcal{F}$ as Bernoulli distribution such that $A(i,j)\sim\mathrm{Bernoulli}(\Omega(i,j))$. Let $p=0.2$ and $\rho$ range in $\{0.1,0.1,\ldots,1\}$. Panels (d)-(f) of Figure 3 show the results. For the task of estimating memberships, all methods enjoy better performances as $\rho$ increases which is consistent with our analysis in Example 2. For the task of determining $K$, all methods have larger accuracy rates when $\rho$ increases.
@@
Experiment 3: Poisson distribution. Set $\mathcal{F}$ as Poisson distribution such that $A(i,j)\sim\mathrm{Poisson}(\Omega(i,j))$. Let $p=0.2$ and $\rho$ range in $\{0.2,0.4,\ldots,2\}$. The results displayed in panels (g)-(i) of Figure 3 are similar to that of Experiment 2 and we omit the analysis here.
@@
Experiment 4: Uniform distribution. Set $\mathcal{F}$ as Uniform distribution such that $A(i,j)\sim\mathrm{Uniform}(0,2\Omega(i,j))$. Let $p=0.2$ and $\rho$ range in $\{2,4,\ldots,20\}$. Panels (j)-(l) of Figure 3 display the results. We observe the following: (i) ScD’s performance is almost unchanged when $\rho$ becomes larger, which is consistent with our analysis in Example 4. (ii) ScD, GeoNMF, SPACL, and Mixed-SCORE perform satisfactorily and similarly because their Hamming errors are lesser than 0.07, their Relative errors are lesser than 0.18, and their ONMs are larger than 0.65. (iii) Our KScD, KSPACL, and KMixed-SCORE correctly determine $K$ for this experiment. It is interesting to see that NB and BHac successfully estimate $K$ when $\rho<10$ while they fail to infer $K$ when $\rho\geq 10$.
@@
Experiment 5: Signed network. For signed network when $\mathbb{P}(A(i,j)=1)=\frac{1+\Omega(i,j)}{2}$ and $\mathbb{P}(A(i,j)=-1)=\frac{1-\Omega(i,j)}{2}$. Let $p=-0.2$ and $\rho$ range in $\{0.1,0.2,\ldots,1\}$. The last three panels of Figure 3 display the results. We see that ScD, GeoNMF, SPACL, and Mixed-SCORE perform similarly and they behave better in estimating memberships when $\rho$ increases, which supports our analysis in Example 5. For determining $K$, the accuracy rates of our methods KScD, KSPACL, and KMixed-SCORE increase when increasing $\rho$ while their two competitors fail to find $K$ for signed networks.
(a) Normal distribution
(b) Normal distribution
(c) Normal distribution
(d) Bernoulli distribution
(e) Bernoulli distribution
(f) Bernoulli distribution
(g) Poisson distribution
(h) Poisson distribution
(i) Poisson distribution
(j) Uniform distribution
(k) Uniform distribution
(l) Uniform distribution
(m) Signed networks
(n) Signed networks
(o) Signed networks
Figure 3: Numerical results of Experiments 1-5.
@@
  • 1.
    For the task of estimating memberships, the results support our analysis in Examples 1-5 that ScD performs differently for different distributions when we change $\rho$. GeoNMF, SPACL, and Mixed-SCORE have competitive performances with ScD and they can estimate memberships for networks generated from DCMMDF.
  • 2.
    For the task of determining the number of communities, our methods enjoy satisfactory performances, which suggests the effectiveness of our overlapping weighted modularity. For comparison, NB and BHac fail to estimate $K$ when $\mathcal{F}$ is Normal distribution, Uniform distribution, and signed network.

### v3

@@ 5 Estimation of the Number of Communities
In this section, we aim to answer Q (3). Algorithm 1 requires a known number of communities $K$. However, for real-world networks, $K$ is usually unknown, which diminishes the usefulness of a community detection approach. Here, to estimate $K$, we introduce overlapping weighted modularity.
@@
$Q^{+}=\frac{1_{m^{+}>0}}{m^{+}}\sum_{k=1}^{K}\sum_{i\in\hat{\mathcal{C}}_{k},j\in\hat{\mathcal{C}}_{k}}\frac{1}{O(i)O(j)}(A^{+}(i,j)-\frac{d^{+}(i)d^{+}(j)}{m^{+}}),$
$Q^{-}=\frac{1_{m^{-}>0}}{m^{-}}\sum_{k=1}^{K}\sum_{i\in\hat{\mathcal{C}}_{k},j\in\hat{\mathcal{C}}_{k}}\frac{1}{O(i)O(j)}(A^{-}(i,j)-\frac{d^{-}(i)d^{-}(j)}{m^{-}}),$
where $1_{m^{+}>0}$ is indicator function such that $1_{m^{+}>0}=1$ if $m^{+}>0$ and 0 otherwise. Similar for $1_{m^{-}>0}$. [68] discovered that social networks often exhibit assortativity, where nodes preferentially connect to others similar to them. Specifically, nodes within the same group tend to have more connections than nodes from different groups [68, 69, 70, 71, 72]. Conversely, technological and biological networks often display dis-assortativity, where nodes in the same group have fewer connections than nodes in different groups. Following [73], our WDCMM model can generate assortative weighted networks by setting the maximum off-diagonal element of $P$ smaller than its minimum diagonal element. Conversely, dis-assortative weighted networks can be modeled by setting the minimum off-diagonal element of $P$ larger than its maximum diagonal element. In this paper, we introduce $\eta_{\mathcal{N}}$ as an index to represent the assortativity of a weighted network $\mathcal{N}$, where $\eta_{\mathcal{N}}=1$ indicates an assortative network, and $\eta_{\mathcal{N}}=-1$ indicates a dis-assortative network. Inspired by the concept of weighted modularity for signed networks [74], we define our overlapping weighted modularity as follows:
$Q_{ovw}=(\frac{m^{+}}{m^{+}+m^{-}}Q^{+}-\frac{m^{-}}{m^{+}+m^{-}}Q^{-})\eta_{\mathcal{N}},$ (13)
@@
where ovwovw means “overlapping weighted”. This measure evaluates the quality of overlapping community partition in both assortative and dis-assortative weighted networks. When $\eta_{\mathcal{N}}=1$ (assortative network), $Q_{ovw}=\frac{m^{+}}{m^{+}+m^{-}}Q^{+}-\frac{m^{-}}{m^{+}+m^{-}}Q^{-}$, and it quantifies the quality of overlapping community partitions in assortative networks. When $\eta_{\mathcal{N}}=-1$ (dis-assortative network), $Q_{ovw}=\frac{m^{-}}{m^{+}+m^{-}}Q^{-}-\frac{m^{+}}{m^{+}+m^{-}}Q^{+}$, and it assesses the quality of overlapping community detection in dis-assortative networks. Similar to the Newman-Girvan modularity [75, 64] and weighted modularity [74], a larger overlapping weighted modularity indicates a better community partition. Our overlapping weighted modularity includes several popular modularity as special cases when $\eta_{N}=1$ (assortative network).
@@
  • When $A$ has both positive and negative entries and $\hat{\Pi}^{0}$ degenerates such that each node only belongs to one community, our overlapping weighted modularity reduces to the modularity for signed networks provided in Equation (17) of [74].
@@
  • When all entries of $A$ are nonnegative (i.e., $m^{-}=0$), our overlapping weighted modularity degenerates to the overlapping modularity provided in Equation (2) of [76].
@@
  • When $\hat{\Pi}^{0}$ degenerates such that each node only belongs to one community and all entries of $A$ are nonnegative, our overlapping weighted modularity degenerates to the famous Newman-Girvan modularity [75, 64].
To infer the number of communities, we adopt the suggestion proposed in [77, 78], that is, we iteratively increase $k$ and choose the one maximizing our overlapping weighted modularity in Equation (13). We let $\hat{\Pi}^{0}$ be obtained from our ScD algorithm and call our method for inferring $K$ as KScD. Given our newly defined overlapping weighted modularity $Q_{ovw}$ for measuring the quality of overlapping community partitions in assortative and dis-assortative weighted networks, it is natural to question its effectiveness. To address this, we will examine the Accuracy rate (defined later) of our KScD method in estimating $K$ for both assortative and dis-assortative weighted networks in the next section. If our KScD (designed by maximizing $Q_{ovw}$) estimates $K$ with a high Accuracy rate, it indicates that our modularity $Q_{ovw}$ is a reliable measure. Numerical results in the subsequent section confirm that our overlapping weighted modularity $Q_{ovw}$ is indeed a valid measure.
@@ 6 Experiments
We conduct extensive experiments to validate our theoretical results, demonstrate the effectiveness of ScD in estimating membership matrix $\Pi$, and demonstrate the effectiveness of our overlapping weighted modularity by showing that KScD can infer $K$ for both assortative and dis-assortative networks generated from WDCMM. The experimental part is implemented with Matlab R2021b on a standard personal computer (Thinkpad X1 Carbon Gen 8).
@@ 6.1 Baseline Methods
For the task of estimating memberships, we select three mixed membership community detection approaches as baseline methods of our WDCMM model.
  • GeoNMF [79] adapts nonnegative matrix factorization to estimate mixed memberships for networks generated from MMSB.
  • SPACL [36] is a spectral algorithm for estimating mixed memberships for networks generated from MMSB.
  • Mixed-SCORE [33] and OCCAM [34] are two spectral methods for estimating mixed memberships for networks generated from DCMM.
@@
GeoNMF, Mixed-SCORE, and OCCAM may fail to output for the case that some elements in the adjacency matrix $A$ are negative. To make them work when $A$ has negative elements, we make all entries of $A$ nonnegative by adding a sufficiently large constant for these three methods.
@@
  • NB and BHac [80] are two methods for inferring the number of communities and they determine $K$ via non-backtracking matrix and Bethe Hessian matrix, respectively.
@@ 6.2 Evaluation metrics
Metrics with ground truth. For this case, we introduce three widely used evaluation metrics: Hamming error [33] and Relative error [36].
@@
  • Overlapping normalized mutual information (ONMI) defined in [81] can also be applied to measure the effectiveness of the proposed methods while comparing with ground truth. Since ONMI works for binary overlapping memberships [34, 35] while $\Pi\in[0,1]^{n\times K}$ and $\hat{\Pi}\in[0,1]^{n\times K}$ are continuous memberships in this paper, to compute ONMI, we use $\Pi^{0}$ computed from $\Pi$ and $\hat{\Pi}^{0}$ computed from $\hat{\Pi}$. ONMI is no larger than 1 and a larger ONMI means a better estimation of the membership matrix $\Pi$.
Metrics without ground truth. For this case, our overlapping weighted modularity in Equation (13) is capable of evaluating the quality of detected communities, and a higher value of overlapping weighted modularity indicates more accurately detected communities.
Accuracy rate. For the task of estimating $K$, we use the Accuracy rate to evaluate the precision of KScD and its competitors, where Accuracy rate is defined as the proportion of times an approach exactly determines $K$ [80].
@@ 6.3 Simulations
(a) $\eta_{\mathcal{N}}=1$
(b) $\eta_{\mathcal{N}}=1$
(c) $\eta_{\mathcal{N}}=1$
(d) $\eta_{\mathcal{N}}=1$
(e) $\eta_{\mathcal{N}}=1$
(f) $\eta_{\mathcal{N}}=-1$
(g) $\eta_{\mathcal{N}}=-1$
(h) $\eta_{\mathcal{N}}=-1$
(i) $\eta_{\mathcal{N}}=-1$
(j) $\eta_{\mathcal{N}}=-1$
Figure 3: Numerical results of Experiment 1.
For computer-generated networks, we set $n=400$ and $K=3$. Let each community have $n_{0}$ number of pure nodes. For pure node $i$, let $\Pi(i,k)=1$ if $1+(k-1)n_{0}\leq i\leq kn_{0}$ for $i\in[Kn_{0}],k\in[K]$. Set $n_{0}=40$. Let all mixed nodes have four different memberships $(0.1,0.1,0.8),(0.1,0.8,0.1),(0.8,0.1,0.1)$, and $(1/3,1/3,1/3)$, each with $\frac{n-3n_{0}}{4}$ number of nodes. Let $P$’s diagonal elements be 1 and non-diagonal elements be $p$. For Normal distribution and signed networks, $p$ can be negative. For other distributions, $p$ should be positive. Let $\theta(i)=\sqrt{\rho}\times(\mathrm{rand}(1)/2+0.5)$, where $\mathrm{rand}(1)$ is a random value ranged in $(0,1)$. In this way, $n,K,P,\Pi$, and $\Theta$ have been set. Then following Steps (a)-(c), we can generate an adjacency matrix $A$ with true membership matrix $\Pi$ for any distribution $\mathcal{F}$ satisfying Equation (6) under our WDCMM model. For the task of community detection, applying ScD (and its competitors) to $A$ with $K$ communities obtains the estimated membership matrix $\hat{\Pi}$, and then we can obtain Hamming error and Relative error. For each parameter setting considered in this section, we report the averaged Hamming error (Relative error, and ONMI) over 50 repetitions for each method. We also report the average running time for each method over the 50 repetitions. For the task of determining the number of communities, applying KScD (and its competitors) to $A$ obtains the estimated number of communities, and then we can obtain the Accuracy rate over the 50 repetitions.
@@
For our synthetic networks, $P,\Pi$, and $\Theta$ can be set arbitrarily as long as Equations (2)-(5) hold and $P$’s non-diagonal elements should be positive or nonnegative or can be negative depending on distribution $\mathcal{F}$ just as analyzed in Examples 1-5. Meanwhile, for simulations, $\mathcal{F}$ can be any distribution since our WDCMM is a distribution-free model and it has no constraint on a specific distribution $\mathcal{F}$ as long as Equation (6) holds. Furthermore, if we consider missing edges, ScD performs better as the number of 1s in $\mathcal{A}$ increases. Therefore, to save space, we do not consider missing edges in the simulations.
Experiment 1: Normal distribution. Set $\mathcal{F}$ as Normal distribution such that $A(i,j)\sim\mathrm{Normal}(\Omega(i,j),\sigma^{2}_{A})$. Let $\sigma^{2}_{A}=2$ and $\rho$ range in $\{0.1,0.2,\ldots,2\}$. For assortative weighted networks (i.e., $\eta_{\mathcal{N}}=1$), we set p=0.2p=-0.2. For dis-assortative weighted networks (i.e., $\eta_{\mathcal{N}}=-1$), we set $p=5$. The results are in Figure 3. We observe the following: (i) ScD performs better as $\rho$ increases, and this is consistent with our theoretical analysis in Example 1. (ii) For assortative networks, all methods perform similarly and they all successfully estimate memberships for large $\rho$. For dis-assortative networks, GeoNMF and OCCAM fail to estimate mixed memberships while the other three methods behave similar and perform better when we increase $\rho$. Meanwhile, SPCAL runs slightly faster than ScD, Mixed-SCORE, and OCCAM while GeoNMF runs slowest. (iii) For the task of inferring the number of communities, our methods KScD, KSPACL, and KMixed-SCORE designed based on our overlapping weighted modularity enjoy similar and satisfactory performances because their accuracy rates increase as $\rho$ increases, which indicates the effectiveness of our overlapping weighted modularity for both assortative and dis-assortative networks. For comparison, NB and BHac fail to estimate $K$ for Normal distribution. (iv) We observe that $\mathrm{ONMI}<0.7$ for all methods (this observation is also found in Experiments 2-5). This is acceptable since ONMI works for binary overlapping memberships and we compute it by using $\Pi^{0}$ and $\hat{\Pi}^{0}$, where $\Pi^{0}$ (and $\hat{\Pi}^{0}$) differs from $\Pi$ (and $\hat{\Pi}$), i.e., using ONMI to measure the difference between $\Pi^{0}$ and $\hat{\Pi}^{0}$ can not fully capture the difference between $\Pi$ and $\hat{\Pi}$.
@@
For the task of estimating memberships, in our experimental studies, we aim at using simulations to verify our theoretical analysis that ScD has different behaviors for different distributions when $\rho$ is changed as analyzed in Examples 1-5. We also want to investigate the performances of GeoNMF, SPACL, Mixed-SCORE, and OCCAM when they are applied to estimate memberships for weighted networks generated from our model WDCMM instead of showing that ScD outperforms its competitors.
(a) $\eta_{\mathcal{N}}=1$
(b) $\eta_{\mathcal{N}}=1$
(c) $\eta_{\mathcal{N}}=1$
(d) $\eta_{\mathcal{N}}=1$
(e) $\eta_{\mathcal{N}}=1$
(f) $\eta_{\mathcal{N}}=-1$
(g) $\eta_{\mathcal{N}}=-1$
(h) $\eta_{\mathcal{N}}=-1$
(i) $\eta_{\mathcal{N}}=-1$
(j) $\eta_{\mathcal{N}}=-1$
Figure 4: Numerical results of Experiment 2.
(a) $\eta_{\mathcal{N}}=1$
(b) $\eta_{\mathcal{N}}=1$
(c) $\eta_{\mathcal{N}}=1$
(d) $\eta_{\mathcal{N}}=1$
(e) $\eta_{\mathcal{N}}=1$
(f) $\eta_{\mathcal{N}}=-1$
(g) $\eta_{\mathcal{N}}=-1$
(h) $\eta_{\mathcal{N}}=-1$
(i) $\eta_{\mathcal{N}}=-1$
(j) $\eta_{\mathcal{N}}=-1$
Figure 5: Numerical results of Experiment 3.
@@
Experiment 2: Bernoulli distribution. Set $\mathcal{F}$ as Bernoulli distribution such that $A(i,j)\sim\mathrm{Bernoulli}(\Omega(i,j))$. For assortative networks (i.e., $\eta_{\mathcal{N}}=1$), we let $p=0.2$ and $\rho$ range in $\{0.1,0.2,\ldots,1\}$. For dis-assortative networks (i.e., $\eta_{\mathcal{N}}=-1$), we let $p=5$ and $\rho$ range in $\{0.02,0.04,\ldots,0.2\}$. Figure 4 shows the results. For the task of estimating memberships, with the exception of GeoNMF and OCCAM, which fail to work for dis-assortative networks, all methods enjoy better performances as $\rho$ increases which is consistent with our analysis in Example 2. Meanwhile, we see that SPCAL and ScD run faster than the other three methods. For the task of determining $K$, all methods have larger accuracy rates when $\rho$ increases in assortative networks. Notably, our KScD, KSPACL, and KMixed-SCORE methods also exhibit high accuracy in estimating $K$, whereas NB and BHac fail to work for dis-assortative networks.
(a) $\eta_{\mathcal{N}}=1$
(b) $\eta_{\mathcal{N}}=1$
(c) $\eta_{\mathcal{N}}=1$
(d) $\eta_{\mathcal{N}}=1$
(e) $\eta_{\mathcal{N}}=1$
(f) $\eta_{\mathcal{N}}=-1$
(g) $\eta_{\mathcal{N}}=-1$
(h) $\eta_{\mathcal{N}}=-1$
(i) $\eta_{\mathcal{N}}=-1$
(j) $\eta_{\mathcal{N}}=-1$
Figure 6: Numerical results of Experiment 4.
@@
Experiment 3: Poisson distribution. Set $\mathcal{F}$ as Poisson distribution such that $A(i,j)\sim\mathrm{Poisson}(\Omega(i,j))$. Let $p=0.2$ for assortative weighted networks (i.e., $\eta_{\mathcal{N}}=1$), $p=5$ for dis-assortative weighted networks (i.e., $\eta_{\mathcal{N}}=-1$), and $\rho$ range in $\{0.2,0.4,\ldots,2\}$. The results displayed in Figure 5 are similar to that of Experiment 2 and we omit the analysis here.
@@
Experiment 4: Uniform distribution. Set $\mathcal{F}$ as Uniform distribution such that $A(i,j)\sim\mathrm{Uniform}(0,2\Omega(i,j))$. Let $p=0.2$ for assortative weighted networks (i.e., $\eta_{\mathcal{N}}=1$), $p=5$ for dis-assortative weighted networks (i.e., $\eta_{\mathcal{N}}=-1$), and $\rho$ range in $\{2,4,\ldots,20\}$. Figure 6 displays the results. We observe the following: (i) ScD’s performance is almost unchanged when $\rho$ becomes larger, which is consistent with our analysis in Example 4. (ii) for assortative networks, all methods perform satisfactorily and similarly because their Hamming errors are lesser than 0.07, their Relative errors are lesser than 0.18, and their ONMIs are larger than 0.65. For dis-assortative networks, ScD, SPACL, and Mixed-SCORE enjoy satisfactory performances while GeoNMF and OCCAM fail to work. (iii) ScD, SPACL, and GeoNMF run faster than Mixed-SCORE and OCCAM in assortative networks while GeoNMF is the slowest in the dis-assortative case. (iv) Our KScD, KSPACL, and KMixed-SCORE correctly determine $K$ for this experiment in both assortative and dis-assortative networks. It is interesting to see that NB and BHac successfully estimate $K$ when $\rho<10$ while they fail to infer $K$ when $\rho\geq 10$ in assortative networks and these two methods fail to determine $K$ for dis-assortative networks.
@@
Experiment 5: Signed network. For signed network when $\mathbb{P}(A(i,j)=1)=\frac{1+\Omega(i,j)}{2}$ and $\mathbb{P}(A(i,j)=-1)=\frac{1-\Omega(i,j)}{2}$. Let p=0.2p=-0.2 and $\rho$ range in $\{0.1,0.2,\ldots,1\}$ for assortative signed networks (i.e., $\eta_{\mathcal{N}}=1$). Let $p=5$ and $\rho$ range in $\{0.02,0.04,\ldots,0.2\}$ for dis-assortative signed networks (i.e., $\eta_{\mathcal{N}}=-1$). Figure 7 displays the results. We see that: (i) ScD exhibits improved performance in estimating memberships as $\rho$ increases, which supports our analysis in Example 5. (ii) All methods, with the exception of GeoNMF and OCCAM, demonstrate similar behavior and improved performance as $\rho$ grows in both assortative and dis-assortative networks. For this experiment, SPACL runs faster than the other four methods while GeoNMF runs slowest. For determining $K$, the accuracy rates of our methods KScD, KSPACL, and KMixed-SCORE increase when increasing $\rho$ while their two competitors fail to find $K$ for signed networks.
(a) $\eta_{\mathcal{N}}=1$
(b) $\eta_{\mathcal{N}}=1$
(c) $\eta_{\mathcal{N}}=1$
(d) $\eta_{\mathcal{N}}=1$
(e) $\eta_{\mathcal{N}}=1$
(f) $\eta_{\mathcal{N}}=-1$
(g) $\eta_{\mathcal{N}}=-1$
(h) $\eta_{\mathcal{N}}=-1$
(i) $\eta_{\mathcal{N}}=-1$
(j) $\eta_{\mathcal{N}}=-1$
Figure 7: Numerical results of Experiment 5.
@@
  • For the task of estimating memberships, the numerical results in Examples 1-5 support our analysis that ScD exhibits varying performance for different edge weight distributions as $\rho$ changes. GeoNMF, SPACL, Mixed-SCORE, and OCCAM show competitive performances with ScD and can estimate memberships for assortative networks generated from the WDCMM model. However, for dis-assortative networks, GeoNMF and OCCAM fail to work, while ScD, SPACL, and Mixed-SCORE demonstrate similar and improved performance as $\rho$ increases.
  • Our methods enjoy satisfactory performance in determining the number of communities in both assortative and dis-assortative weighted networks, highlighting the effectiveness of our overlapping weighted modularity $Q_{ovw}$. Thus, a higher $Q_{ovw}$ value indeed indicates a better community partition. Conversely, NB and BHac perform well only in assortative networks generated from Bernoulli and Poisson distributions. They fail to estimate $K$ for assortative networks derived from Normal and Uniform distributions, as well as signed networks. Moreover, they entirely fail to infer $K$ for dis-assortative networks.
Remark 11.
This discrepancy in performance of different methods stems from a fundamental methodological difference. The superiority of ScD, SPACL, and Mixed-SCORE stems from their distribution-free spectral approach, which utilizes the eigenvectors corresponding to the top $K$ eigenvalues in magnitude of the adjacency matrix $A$. This strategy impartially captures the strongest structural signals, whether they manifest as large positive or large negative eigenvalues—the latter being a characteristic of dis-assortative networks. In fact, in the proof of Theorem 1, we only require the connectivity matrix $P$ to satisfy Equation (4) without limiting it to be specifically designed for assortative or dis-assortative networks. Conversely, GeoNMF, grounded in non-negative matrix factorization, is fundamentally constrained by its requirement for non-negative inputs and factor matrices. This constraint renders it incapable of representing the core dis-assortative structural pattern where connections between different communities are stronger than those within communities. Similarly, OCCAM relies solely on the largest positive eigenvalues, thereby disregarding the dominant negative spectral information that is critical in dis-assortative networks. The flexibility of the spectral approach employed by ScD and its peers is thus the key to their universal applicability across both assortative and dis-assortative networks.